OpenStudy (anonymous):

√14 is between which two whole numbers? A. 3 and 4 B. 4 and 5 <--- C. 6 and 7 D. 7 and 8 3x3=9 4x4=16 √14 fits right in between 9 and 16? @campbell_st @Embryo

OpenStudy (nikosis01):

b seem right

OpenStudy (anonymous):

You need to look at it like this, you're asked to find some number that sqrt{14} is greater than, but also another number that it's less than\[a<\sqrt{14}<b\]if you square all of the terms you would get\[a^2<14<b^2\]the you just look at it and see what number squared is less than 14, and what number squared is greater than 14

OpenStudy (anonymous):

WRONG

OpenStudy (campbell_st):

well if you think 3^2 = 9 and 4^2 = 16 \[(\sqrt{14})^2 = 14\] so where does it lie...?

OpenStudy (nikosis01):

wats wrong

OpenStudy (anonymous):

3 and 4

OpenStudy (anonymous):

Yup, it's A...

OpenStudy (nikosis01):

It seem like b but yea its a

OpenStudy (anonymous):

Oh sorry! :) I did mean A. :) LOL, My explanation doesn't even fit with B. :) Thanks @nikosis01 @iambatman @zharper12345678910 @campbell_st @Embryo

OpenStudy (anonymous):

your wellcome

OpenStudy (nikosis01):

ur welcome CrazyCountryGirl